Kevin Geiss

28248 N. 90th Ln Peoria, AZ 85383 +1 (602) 881-9265 kevin.geiss@gmail.com

astrophoenix.com

Twenty-five years professional experience developing large-scale software projects in various UNIX-style environments. Solid knowledge of algorithms, complexity, and computer science theory

Skills

Languages Java, Spring, Perl, C++, C, JavaScript, Swift, TCP, TEX, HTML, SQL, Ada, Pascal, Fortran 90, Assembly, Bash or Zsh scripting

Version Control Git, Git Submodules, Git-lfs, Git-annex, Mercurial, Subversion, CVS, RCS, Clear Case, and SCCS

IDEs/tools Linux, Apache, Docker, Kubernetes, Cloudstack, IntelliJ, Eclipse, Xcode, Maven, Vim, GNU Make, GNU Screen, Zsh

Experience

{2020– \atop Present} Senior Software Development Engineer (SDE3), Amazon, Tempe, AZ.

2002–2020 Principal Engineer I, Ticketmaster, Scottsdale, AZ.

Implemented web and network applications on Apache and Linux. Ported network applications from Windows and VAX to Linux.

Containerized Perl and C++ products to run on Kubernetes Deployed with Helm and Tiller

Created a Gatling Capacity Test suite for the Company-wide NFL initiative Tuned products from four teams, which form the B2B Secondary Market stack Adjusted Apache configurations, VIP setup, and system configurations Increased the existing system’s capacity 100-fold, to handle sustained 320 RPS

Mentored teams on technologies, processes, and platforms Docker, Kubernetes, Linux Java and Perl Unit testing and Test-Driven Development Code reviews Integrating with other services Git (feature branches, merging, rebasing, submodules, git-lfs, git-annex)

Created a public-facing webapp to render Tickets on mobile devices Implemented asynchronous calls in Java 8 to backend services Used mustache templates for rendering the frontend Created a B2B point-of-sale REST API using Java, Spring, CXF on Tomcat Integrated with other services using jaxrs:client proxies 100% unit test coverage of all application code (Java and JavaScript) Responsible for the platform engineering: Continuous integration - Jenkins build pipeline, Rundeck for deployments - Provisioning and configuration of the virtual servers Wrote perl scripts to document releases, and to parse log files for debugging

Wrote JavaScript code to generate a public/private keypair and make a Certificate Signing Request, replacing the deprecated “keygen” browser tag

Implemented SSL Certificate Pinning in an iOS app (Swift and Objective-C)

Shepherded multiple products through the PCI certification process Documented the protections for Sensitive Cardholder Data Presented the documentation to the PCI Auditors through live interviews

Spearheaded unit testing Swift code Picked MockFive after evaluating Swift mocking frameworks for unit tests

Refactored a B2B web app (Java, JBoss and Seam) to increase performance Adopted standard Seam design patterns to speed up the database interaction Changed runtime complexity from O(n^{2}) throughout the app to O(n) in some cases, O(n logn) in others, by changing data structures

Created a B2B web service for Resale/Secondary Market ticket exchange and trans- fer (ModPerl, Apache) Designed and implemented an Actions framework to encapsulate business logic into reusable pieces which became the basis for our other ModPerl web services. Refactored multiple B2C apps into the class hierarchy to maximize code reuse and increase their functionality

Created a B2B service to provide live onsale data to clients (C++) One server handled hundreds of simultaneous requests Implemented a priority queue using a heap to ensure data returned to clients at specified intervals Used the select system call to manage the pool of sockets

Ported a C++ request queueing engine from Windows to Linux which services tens of thousands of simultaneous requests

Modified a Windows-only C++ service and utility framework so it is now cross plat- form, running on both Linux and Windows. TMLib is used in all other Ticketmaster C++ applications.

2013–2018 Database Architect, Land of Frankincense Archaeological Project.

ualr.edu/anthropology/archaeology/land-of-frankincense Designed and implemented Filemaker database to store all Archaeological data from the dig

Created over 100 tables with one-to-many and many-to-many relationships between multiple

sets of tables. Collaborated with the UI designer to implement Forms for inputting data in the field,

leveraging the table relationships to pre-fill generated data where possible.

2000–2002 Software Engineer, VoiceViewer Technologies, Inc., and Gordon Minns & Associates, Scottsdale, AZ.

The VoiceViewer is a cell phone with integrated microdisplay; resolution of 800x600 pixels with 24-bit color

Implemented Linux device driver, console, and Linux kernel frame buffer in C to abstract

the microdisplay for application use (e.g., XFree86) Modified XFree86 for compatibility with the microdisplay Implemented image processing utilities in C++ to pre-process bitmaps for use on the mi-

crodisplay

2000–2001 Software Engineer, Sensory Science Corporation, Scottsdale, AZ.

The EDGE Media Appliance (convergence entertainment device on a Linux platform)

Detailed software design and C++ implementation of the media appliance software Created the software build system of Makefiles which compile the software over a hierarchy

of directories Designed, implemented, and administered the Linux development network, including

nightly backups, revision control software (CVS), firewall, NFS File server, NIS (yp), and Domain Name Server (DNS)

1998–2000 Senior Staff Analyst, Compuware Corporation, Phoenix, AZ.

Under contract to American Express Corporation, International Risk Management Automation

Member of the Infrastructure Team. Responsible for creating C++ tools for the development and maintenance of Virtual Analysts (VAs) which automate human analysts’ tasks on American Express mainframes.

Implemented utilities in C++ and Perl to secure VA login passwords using strong cryptog-

raphy Designed and implemented a C++ application which monthly changes the passwords of

1500 VA IDs on 22 heterogenous mainframe systems. The application replaces a nine-hour manual procedure. Supporting additional mainframe systems requires only linking in new code; existing code needs no recompilation (avatar factory pattern).

Examined and repaired legacy C++ code for Y2K problems

1996–1998 Software Engineer, Lockheed Martin Tactical Defense Systems, Goodyear, AZ.

Helped develop a large-scale data processing application

Participated in high-level and detailed software design of the C++ application. Called upon

to make critical design decisions Researched and Prototyped calls to commercial and vendor libraries to discover their best

application Created the software build system of Makefiles which compile the software over a hierarchy

of directories

1995–1996 Systems Programmer, Lunar & Planetary Lab, University of Arizona, Tucson, AZ.

The NASA Imager for Mars Pathfinder (IMP) project created a CCD stereo imaging device which traveled with the NASA Mars Pathfinder mission to the Martian surface.

Developed real-time, embedded software in C to control the imager and interface with the

lander. The system on the lander was VxWorks on a VME platform. Created OpenGL application on Silicon Graphics IRIX to display the stereo images in 3D. Wrote software to collate data from dozens of sensors used in the temperature cycling tests

1995 Graduate Teaching Assistant, University of Arizona Computer Science Department, Tucson, AZ.

Assisted with CSC432 (Data Structures and Algorithms)

Professional Training

Ticketmaster  Lockheed Martin

AWS Certified Solution Architect Asso-

ciate BootCamp Dynamic Leadership Training DevOps Training Swift/iOS Training Cucumber BDD Training Agile Training Perl Training C++ Training Digital Signal Processing

Free Software Contributions

2006 ABClock, www.let.rug.nl/~kleiweg/abclock/, Version 1.0d.

Added support for X11 application dock withdrawn state

2005 ProcPS Free, astrophoenix.com/gitlist/free, Version 3.1.9.

Ported the “free” command-line memory display utility from Linux’s procps to Mac OS X

2003 LavaPS, www.isi.edu/~johnh/SOFTWARE/LAVAPS/, Version 2.0.

Contributed Mac OS X support

2002–2004 Gentoo Linux PPC on Apple iBook, astrophoenix.com/ppc, Linux 2.6.7.

Maintained information and configurations for installing Gentoo Linux on an Apple iBook

Awards

Nov. 2015 Mark 1 Award, Ticketmaster, Scottsdale, AZ.

Nominated by peers as the most valuable contributor to the current iteration.

“I am nominating Kevin G for his contribution towards Identity Adapter integration in

WPS and Ticketing Service which unblocked all the three teams to move forward.” “I would like to nominate Kevin Geiss for the Mark 1 Award for the work he has been doing

for the Identity Adapter and always being available to help/debug WPS issues as they arise. Plus he helps everyone on his team get their work done.”

“Kevin Geiss gets my nomination for this iteration for his dedication in getting the SalesDeck Application up in Pre-prod. He went above and beyond to help others and took passion in his own work to resolve all the issues that arose in the Pre-prod effort. Thanks Kevin! We would not have gotten this far without your dedication and hard work!”

“I would like to nominate Kevin for the Mark 1 Award. His work over the last few weeks to get the identity adapter integrated with our dependencies has finally paid off! We recognized this would be a major challenge right from the get-go, and Kevin stepped up to solve this problem.”

“To get SalesDeck working in Portal, Kevin Geiss took on the tasks of verifying Identity Adapter in all environments and got WPS and Ticketing to work with Identity Adapter. Kevin is a great team member. He leads by example. He is very easy to work with and always ready to help his teammates.”

Education

1994–1996 Master of Science, University of Arizona, Tucson, Arizona, 3.4 .

Computer Science

1990–1994 Bachelor of Science, with Distinction in Computer Science, Missouri State University, Springfield, Missouri, 3.63 .

Computer Science & Mathematics Double Major; Physics Minor